# Geodesic Distance Between Graphs: A Spectral Metric for Assessing the Stability of Graph Neural Networks

### v2

@@ 4.1 Phase 1: Spectral Graph Matching for Finding Node Correspondence
Computing the GGD metric between two input graphs requires solving a graph-matching problem in advance. Without knowing the node-to-node correspondence that can be achieved through a graph-matching step, the distance between modified Laplacian matrices may be significantly higher than the minimum possible distance. In this work, we aim to find the infimum between two SPD matrices on the Riemannian manifold formed by modified Laplacian matrices, which can be accomplished through a graph-matching phase. Graph matching techniques can be used to establish node-to-node correspondence by seeking a bijection between node sets to maximize the alignment of edge sets (Livi & Rizzi, 2013; Emmert-Streib et al., 2016; Caetano et al., 2009). This combinatorial optimization problem can be cast into a Quadratic Assignment Problem, which is NP-hard to solve or approximate (Fan et al., 2020; Wang et al., 2020).
Figure 1: Graphs with simple perturbations.
In this study, we will exploit a spectral graph matching method called GRAMPA (GRAph Matching by Pairwise eigen-Alignments) (Fan et al., 2020) to find the approximate node correspondence between two graphs. GRAMPA starts with comparing the eigenvectors of the adjacency matrices of the input graphs. Instead of comparing only the eigenvectors responding to the largest eigenvalues, it considers all pairs of eigenvectors/eigenvalues to generate a similarity matrix. This similarity matrix can be constructed by summing up the outer products of eigenvector pairs, weighted by a Cauchy kernel (Fan et al., 2020). Subsequently, a rounding procedure will be performed to determine the optimal match between nodes employing the similarity matrix.
@@
$\widehat{X}=\sum_{i,j=1}^{n}w\left(\zeta_{i},\mu_{j}\right)\cdot u_{i}u_{i}^{\top}\mathbf{J}v_{j}v_{j}^{\top},\text{where }w(x,y)=\frac{1}{(x-y)^{2}+\eta^{2}}.$ (6)
@@
which can be efficiently solved using the Hungarian algorithm (Fan et al., 2020). However, one simpler rounding procedure was advocated in (Fan et al., 2020) with theoretical results supporting the rounding procedure, which is given by the following equation:
@@
here the permutation estimate matrix is constructed by selecting the largest index from each row. While LAP provides optimal matching, its computational complexity can become expensive for very large graphs. By carefully choosing $\eta$, the same match recovery holds if rounding is performed using equation 7 instead of solving the LAP in equation 8 (Fan et al., 2020).
Lemma 4.1 (Graph Matching Recovery).
Given symmetric matrices $A_{1}$, $A_{2}$ and $Z$ from the Gaussian Wigner model, where $A_{2\pi^{*}}=A_{1}+\sigma Z$, there exist constants $c,c^{\prime}>0$ such that if $1/n^{0.1}\leq\eta\leq c/\log n$ and $\sigma\leq c^{\prime}\eta$, then with probability at least $1-n^{-4}$, GRAMPA Algorithm correctly recovers the permutation matrix $\pi^{*}$ from the Similarity matrix $\widehat{X}$ (Fan et al., 2020). Its proof can be found in the supporting documents A.5.
$\textnormal{Best Match to }A_{2}=\widehat{\pi}A_{1}\widehat{\pi}^{\top},\quad\textnormal{Best Match to }A_{1}=\widehat{\pi}^{\top}A_{2}\widehat{\pi}.$ (9)
In practice, the graph matching performance is not too sensitive to the choice of tuning parameter $\eta$. For small-sized graphs, such as the MUTAG dataset(Morris et al., 2020), setting $\eta=0.5$ yields satisfactory results in matching. In Appendix A.9, the effect of $\eta$ for computing GGDs has been comprehensively analyzed.
@@ 4.2 Phase 2: Computing Geodesic Distances Between Graph Laplacians
The GGD metric can be formally defined as the infimum length of geodesics connecting two data points in the Riemannian manifold formed by the cone of the modified graph Laplacian matrices (Lim et al., 2019). This distance metric can be imagined as a matrix representation of the geometric distance $|\log(a/b)|$ between two positive numbers $a,b$ (Bonnabel & Sepulchre, 2010; Shamai & Kimmel, 2017; Owen & Provan, 2010).
@@
The above GGD formulation for computing distances between SPD matrices is based on an Affine-Invariant Riemannian Metric (AIRM) (Lim et al., 2019), while another well-known metric, the Log-Euclidean Riemannian Metric (LERM) (Ilea et al., 2018; Thanwerdas & Pennec, 2023; Chen et al., 2024) is also discussed in Appendix A.7 of the supplementary section.
@@ 4.3 Connection between GGD and Graph Structural Mismatches
Consider two graphs, $G_{1}$ and $G_{2}$, that have the same node set $V$, with a known correspondence between their nodes. Let $L_{1}$ and $L_{2}$ be the Laplacian matrices of these graphs, respectively. Suppose we take a subset of nodes, denoted by $S$ and its complement, $S^{\prime}$. We assign the value $1$ to the nodes in $S$ and the value 00 to those in $S^{\prime}$. This defines the set $S$ as:
@@
Figure 2: The cut mismatch (for the node set $S$) between two simple graphs is $\frac{6}{2}=3$.
For graph $G_{1}$, the cut for the node subset $S$ (which is the number of edges that cross between $S$ and $S^{\prime}$) can be computed as:
$\text{cut}_{G_{1}}(S,S^{\prime})=x^{T}L_{1}x.$
As shown in Figure 2, for node subset $S$ six edges cross between $S$ and $S^{\prime}$ in graph $G_{1}$, whereas only two edges cross in graph $G_{2}$. This difference in edge counts between the two graphs is referred to as a cut mismatch. The relationship between this cut mismatch and the generalized eigenvalue problem for the matrix pair $(L_{1},L_{2})$ can be formalized using the Generalized Courant-Fischer Minimax Theorem (Golub & Van Loan, 2013; Feng, 2020).
@@
This theorem provides a method to bound the maximum cut mismatch between two graphs by calculating the largest generalized eigenvalue. Specifically, we can use the following optimization problem to compute the dominant eigenvalue $\lambda_{max}$ (Feng, 2020):
$\lambda_{\max}=\max_{\begin{subarray}{c}|x|\neq 0\\ x^{\top}\mathbb{1}=0\end{subarray}}\frac{x^{\top}L_{1}x}{x^{\top}L_{2}x}\geq\max_{\begin{subarray}{c}|x|\neq 0\\ x(v)\in\{0,1\}\end{subarray}}\frac{x^{\top}L_{1}x}{x^{\top}L_{2}x}=\max\frac{cut_{G_{1}}(S,S^{\prime})}{cut_{G_{2}}(S,S^{\prime})}.$ (12)
From equation (12), we can see that the dominant generalized eigenvalue $\lambda_{max}$ corresponds to the most significant cut mismatch between $G_{1}$ and $G_{2}$. In particular, $\lambda_{1}=\lambda_{max}$ sets an upper bound on the cut mismatch between $G_{1}$ and $G_{2}$, while $\lambda_{n}=\lambda_{min}$ defines the upper bound of the mismatch in the reverse direction, between $G_{2}$ and $G_{1}$.
@@ 5 Computing GGDs for Graphs with Different Sizes
To calculate geodesic distances between SPD matrices of different sizes, prior studies have proposed a submatrix adaptation method (Lim et al., 2019). In this approach, a principle submatrix with the same size as the smaller matrix is obtained from the larger matrix (Ye & Lim, 2016), and subsequently used to calculate the GGD. Furthermore, this method can be extended to project the smaller matrix into a larger one with the same size as the larger matrix (Lim et al., 2019). While these methods are efficient for handling SPD matrices, for our application taking the submatrix of the modified Laplacian can lose important nodes/edges, compromising critical graph structural properties.
@@
In this work, we will leverage spectral graph coarsening to address the issue. Spectral graph coarsening is a widely adopted process (Loukas, 2019; Aghdaei & Feng, 2022) for reducing graph sizes while preserving key spectral (structural) properties, such as the Laplacian eigenvalues/eigenvectors. Recent spectral graph coarsening methods aim to decompose an input graph into many distinct node clusters, so that a reduced graph can be formed by treating each node cluster as a new node, with a goal of assuring that the reduced graph will approximately retain the original graph’s structure (Loukas, 2019; Han et al., 2024; Aghdaei & Feng, 2022). Therefore, when computing GGDs for graphs of different sizes, we can first adopt spectral graph coarsening to transform the bigger graph into a smaller one, so that our framework in Section 4 can be subsequently utilized. However, existing state-of-the-art graph coarsening methods do not allow us to precisely control the size of the reduced graphs.
@@ 5.1 Our Approach: Spectral Graph Coarsening by Effective Resistances
In this work, we introduce a spectral graph coarsening method using effective-resistance clustering (Aghdaei & Feng, 2022). Our approach starts with estimating the effective resistances of all edges in the original graph. We can also incorporate the difference between node features (if available) as an additional parameter. In the graph coarsening phase, our method will rank edges according to their resistance distances and only the top few edges with the smallest resistances will be coarsened into new nodes. This approach enables precise control over the size of the reduced graphs while preserving crucial structural properties, such as the eigenvalues/eigenvectors of the adjacency matrices, which are essential for the subsequent spectral graph matching step (Phase 1 in Section 4.1).
Consider a connected, weighted, undirected graph $G=(V,E,w)$ with $|V|=n$. The effective resistance between nodes $(p,q)\in V$ plays a crucial role in various graph analysis tasks including spectral sparsification algorithms (Spielman & Teng, 2011). The effective resistance distances can be accurately computed using the equation:
@@
where ${b_{p}}\in\mathbb{R}^{n}$ denote the standard basis vector with all zero entries except for the $p$-th entry being $1$, and ${b_{pq}}=b_{p}-b_{q}$. $u_{i}\in\mathbb{R}^{n}$ for $i=1,\dots,n$ denote the unit-length, mutually-orthogonal eigenvectors corresponding to Laplacian eigenvalues $\sigma_{i}$ for $i=1,\dots,n$. A brief background on effective resistance is provided in Appendix A.3.
Scalable estimation of effective resistances
To address the computational complexity associated with directly computing eigenvalues and eigenvectors required for estimating edge effective resistances, we leverage a scalable framework for approximating the eigenvectors of the graph Laplacian matrix using the Krylov subspace (Saad, 2011). Let $A$ denote the adjacency matrix of a graph $G$, consider its order-$m$ Krylov subspace $\mathbf{K}_{m}(A,x)$ that is a vector space spanned by the vectors computed through power iterations $x,Ax,A^{2}x,\ldots,A^{m-1}x$ (Liesen & Strakoš, 2012). By enforcing orthogonality among the above vectors in the Krylov subspace, a new set of mutually orthogonal vectors of unit lengths can be constructed for approximating the original Laplacian eigenvectors in 13, which are denoted as $\tilde{u}_{1}$, $\tilde{u}_{2}$, $\ldots$, $\tilde{u}_{m}$. To estimate the effective resistance between two nodes $p$ and $q$, we can exploit the approximated eigenvectors:
$R_{eff}(p,q)\approx\sum\limits_{i=1}^{m}\frac{(\tilde{u}^{\top}_{i}b_{pq})^{2}}{\tilde{u}^{\top}_{i}L\tilde{u}_{i}}+\alpha\|f_{p}-f_{q}\|,$ (14)
Graph coarsening with node features
$R^{*}_{eff}(p,q)=R_{eff}(p,q)+\alpha\|f_{p}-f_{q}\|,$ (15)
where $f_{p}$ and $f_{q}$ are node feature vectors of nodes $p$ and $q$, respectively, while $\alpha$ is a weighting factor that determines the effect of node feature information in the graph coarsening process.
@@ 7.1 Application of GGDs in GNN Stability Analysis
To analyze the stability of GNN models (Bronstein et al., 2017; Garg et al., 2020; Duvenaud et al., 2015), we conducted multiple experiments with the GGD and TMD metrics. GNNs typically operate by a message-passing mechanism (Gilmer et al., 2017), where at each layer, nodes send their feature representations to their neighbors. The feature representation of each node is initialized to its original features and is updated by repeatedly aggregating incoming messages from neighbors. In our experiment, we relate GGD to the Graph Isomorphism Networks (GIN) (Xu et al., 2019), one of the most widely applied and powerful GNNs, utilizing the MUTAG dataset (Morris et al., 2020) as our reference graph dataset. The objective is to analyze the relationship between the input distance $GGD(G_{1},G_{2})$ and the distance between the output GIN vectors, $\|h(G_{1})-h(G_{2})\|$ for randomly selected pairs of graphs.
We employed a three-layer GIN network as described in (Xu et al., 2019). This network uses GIN convolutional layers to update tensors of nodes based on their neighboring nodes and then aggregates those outputs in a vector representation, followed by linear layers for classification tasks. Thus it outputs a single vector $h(G)$ for the entire graph $G$. The result is illustrated in Figure 3.
We observe a strong correlation between GGD and the output distance, as indicated by a high Pearson correlation coefficient. This finding implies the effectiveness of the proposed GGD metric for analyzing the stability of GNN models (Chuang & Jegelka, 2022). To compare GGD with existing metrics, we repeat this experiment using TMD without considering node attributes (features). As shown in Figure 3, GGD demonstrates a better correlation with GIN outputs than the TMD metric across different levels. These findings indicate that when dealing with graphs without node features, GGD should be adopted for the stability analysis of graph learning models.

### v3

@@ 4.1 Phase 1: Spectral Graph Matching for Finding Node Correspondence
Computing the GGD metric between two input graphs requires solving a graph-matching problem in advance., to ensure the minimum possible distance between modified Laplacian matrices. In this work, we aim to find the infimum between two SPD matrices on the Riemannian manifold, which can be accomplished through a graph-matching phase. Graph matching techniques can be used to establish node-to-node correspondence by seeking a bijection between node sets to maximize the alignment of edge sets (livi2013graph, ; emmert2016fifty, ; caetano2009learning, ). This combinatorial optimization problem can be cast into a Quadratic Assignment Problem, which is NP-hard to solve or approximate (fan2020spectral, ; wang2020learning, ).
Figure 2: Graphs with simple perturbations
In this study, we will exploit a spectral graph matching method called GRAMPA (GRAph Matching by Pairwise eigen-Alignments) (fan2020spectral, ) to find the approximate node correspondence between two graphs. GRAMPA starts with comparing the eigenvectors of the adjacency matrices of the input graphs. Instead of comparing only the eigenvectors corresponding to the largest eigenvalues, it considers all pairs of eigenvectors/eigenvalues to generate a similarity matrix. This similarity matrix can be constructed by summing up the outer products of eigenvector pairs, weighted by a Cauchy kernel (fan2020spectral, ). Subsequently, a rounding procedure will be performed to determine the optimal match between nodes employing the similarity matrix.
@@
$\widehat{X}=\sum_{i,j=1}^{n}w\left(\zeta_{i},\mu_{j}\right)\cdot u_{i}u_{i}^{\top}\mathbf{J}v_{j}v_{j}^{\top},\quad w(x,y)=\frac{1}{(x-y)^{2}+\eta^{2}}.$ (6)
@@
which can be efficiently solved using the Hungarian algorithm (fan2020spectral, ). However, one simpler rounding procedure was advocated in (fan2020spectral, ) with theoretical results supporting the rounding procedure, which is given by the following equation:
$\widehat{\pi}(i)=\underset{j}{\operatorname{argmax}}\widehat{X}_{ij},$ (8)
here the permutation estimate matrix is constructed by selecting the largest index from each row. While LAP provides optimal matching, its computational complexity can become expensive for very large graphs. By carefully choosing $\eta$, the same match recovery holds if rounding is performed using equation 7 instead of solving the LAP in equation 8 (fan2020spectral, ).
Lemma 4.2 (Graph Matching Recovery).
Given symmetric matrices $A_{1}$, $A_{2}$ and $Z$ from the Gaussian Wigner model, where $A_{2\pi^{*}}=A_{1}+\sigma Z$, there exist constants $c,c^{\prime}>0$ such that if $1/n^{0.1}\leq\eta\leq c/\log n$ and $\sigma\leq c^{\prime}\eta$, then with probability at least $1-n^{-4}$, GRAMPA Algorithm correctly recovers the permutation matrix $\pi^{*}$ from the Similarity matrix $\widehat{X}$ (fan2020spectral, ). Its proof can be found in the supporting documents A.7.
@@
In practice, the graph matching performance is not too sensitive to the choice of tuning parameter $\eta$. For small-sized graphs, such as the MUTAG dataset(morris2020tudataset, ), setting $\eta=0.5$ yields satisfactory results in matching. In A.15, the effect of $\eta$ for computing GGDs has been comprehensively analyzed.
@@ 4.2 Phase 2: Computing Geodesic Distances Between Graph Laplacians
The GGD metric can be formally defined as the infimum length of geodesics connecting two data points in the Riemannian manifold formed by the cone of the modified graph Laplacian matrices (lim2019geometric, ). This distance metric can be imagined as a matrix representation of the geometric distance $|\log(a/b)|$ between two positive numbers $a,b$ (bonnabel2010riemannian, ; shamai2017geodesic, ; owen2010fast, ).
@@
The above GGD formulation for computing distances between SPD matrices is based on an Affine-Invariant Riemannian Metric (AIRM) (lim2019geometric, ), while another well-known metric, the Log-Euclidean Riemannian Metric (LERM) (ilea2018covariance, ; thanwerdas2023n, ; chen2024adaptive, ) is also discussed in Appendix A.13.
@@ 4.3 Connection between GGD and Graph Structural Mismatches
Consider two graphs, $G_{1}$ and $G_{2}$, that have the same node set $V$, with a known correspondence between their nodes. Let $L_{1}$ and $L_{2}$ be the Laplacian matrices of these graphs, respectively. Suppose we take a subset of nodes, denoted by $S$ and its complement, $S^{\prime}$. We assign the value $1$ to the nodes in $S$ and the value 0 to those in $S^{\prime}$. This defines the set $S$ as:
@@
For graph $G_{1}$, the cut for the node subset $S$ (which is the number of edges that cross between $S$ and $S^{\prime}$) can be computed as: $\text{cut}_{G_{1}}(S,S^{\prime})=x^{T}L_{1}x.$
Figure 3: The cut mismatch (for the node set $S$) between two graphs is $\frac{6}{2}=3$
As shown in Figure 3, for a node subset $S$, six edges cross between $S$ and $S^{\prime}$ in graph $G_{1}$, whereas two edges cross in graph $G_{2}$. This ratio of edge counts in the two graphs is referred as a cut mismatch. The relationship between this cut mismatch and the generalized eigenvalue problem for the matrix pair $(L_{1},L_{2})$ can be formalized using the Generalized Courant-Fischer Minimax Theorem (golub2013matrix, ; feng2020grass, ).
@@
This theorem provides a method for bounding the maximum cut mismatch between two graphs by calculating the largest generalized eigenvalue. Specifically, we can use the following optimization problem to compute the dominant eigenvalue $\lambda_{max}$ (feng2020grass, ):
$\lambda_{\max}=\max_{\begin{subarray}{c}|x|\neq 0\\ x^{\top}\mathbb{1}=0\end{subarray}}\frac{x^{\top}L_{1}x}{x^{\top}L_{2}x}\geq\max_{\begin{subarray}{c}|x|\neq 0\\ x(v)\in\{0,1\}\end{subarray}}\frac{x^{\top}L_{1}x}{x^{\top}L_{2}x}=\max\frac{\mathrm{cut}_{G_{1}}(S,S^{\prime})}{\mathrm{cut}_{G_{2}}(S,S^{\prime})}$ (12)
From equation (12), we can see that the dominant generalized eigenvalue $\lambda_{max}$ corresponds to the most significant cut mismatch between $G_{1}$ and $G_{2}$. In particular, $\lambda_{1}=\lambda_{max}$ sets an upper bound on the cut mismatch between $G_{1}$ and $G_{2}$, while $\lambda_{n}=\lambda_{min}$ defines the upper bound of the mismatch in the reverse direction, between $G_{2}$ and $G_{1}$. Appendix A.8 illustrates this relationship with practical examples. Additionally, we illustrate the relationship between the approximate GGD values with extreme eigenvalues, compared with the accurate GGD values in Appendix A.11.
@@ 5 Computing GGDs for Graphs with Different Sizes
To calculate geodesic distances between SPD matrices of different sizes, prior studies have proposed a submatrix adaptation method (lim2019geometric, ). In this approach, a principle submatrix with the same size as the smaller matrix is obtained from the larger matrix (ye2016schubert, ), and subsequently used to calculate the GGD. Furthermore, this method can be extended to project the smaller matrix into a larger one with the same size as the larger matrix (lim2019geometric, ). While these methods are efficient for handling SPD matrices, for our application taking the submatrix of the modified Laplacian can lose important nodes/edges, compromising critical graph structural properties.
@@
Spectral graph coarsening is a widely adopted process (loukas2019graph, ; aghdaei2022hyperef, ) for reducing graph sizes while preserving key spectral (structural) properties, such as the Laplacian eigenvalues/eigenvectors. Recent spectral graph coarsening methods aim to decompose an input graph into many distinct node clusters, so that a reduced graph can be formed by treating each node cluster as a new node, with a goal of assuring that the reduced graph will approximately retain the original graph’s structure (loukas2019graph, ; han2024topology, ; aghdaei2022hyperef, ). Therefore, when computing GGDs for graphs of different sizes, we can first adopt spectral graph coarsening to transform the bigger graph into a smaller one, so that our framework in Section 4 can be subsequently utilized. However, existing state-of-the-art graph coarsening methods do not allow us to precisely control the size of the reduced graphs.
@@ 5.1 Our Approach: Spectral Graph Coarsening by Effective Resistances
In this work, we exploit a spectral graph coarsening method leveraging effective-resistance clustering (aghdaei2022hyperef, ), specifically designed to reduce graph size while preserving key spectral characteristics. Unlike prior work (lim2019geometric, ), which addresses dimension mismatch directly in SPD space, our coarsening method operates at the structural level before transformation into SPD matrices, making it more suitable for computing distances between graphs with unequal sizes. Our approach begins by estimating the effective resistances of all edges in the original graph. If node features are available, we also incorporate feature differences as an additional parameter. During coarsening, edges are ranked by resistance distance, and only the top few edges with the smallest effective resistances are merged into new nodes. This strategy enables precise control over the size of the reduced graph while preserving crucial structural properties, such as the eigenvalues and eigenvectors of the adjacency matrices—essential for the subsequent spectral graph matching step (Phase 1 in Section 4.1).
Consider a connected, weighted, undirected graph $G=(V,E,w)$ with $|V|=n$. The effective resistance between nodes $(p,q)\in V$ plays a crucial role in various graph analysis tasks including spectral sparsification algorithms (spielman2011spectral, ). The effective resistance distances can be accurately computed using the equation:
@@
where ${b_{p}}\in\mathbb{R}^{n}$ denote the standard basis vector with all zero entries except for the $p$-th entry being $1$, and ${b_{pq}}=b_{p}-b_{q}$. $u_{i}\in\mathbb{R}^{n}$ for $i=1,\dots,n$ denote the unit-length, mutually-orthogonal eigenvectors corresponding to Laplacian eigenvalues $\sigma_{i}$ for $i=1,\dots,n$. Background on effective resistance is presented in Appendix A.4, with its estimation detailed in Appendix A.5.
@@ 7.1 Application of GGDs in GNN Stability Analysis
To analyze the stability of GNN models (bronstein2017geometric, ; garg2020generalization, ; duvenaud2015convolutional, ), we conducted multiple experiments with the GGD and TMD metrics. GNNs typically operate by a message-passing mechanism (gilmer2017neural, ), where at each layer, nodes send their feature representations to their neighbors. The feature representation of each node is initialized to its original features and is updated by repeatedly aggregating incoming messages from neighbors. In our experiment, we relate GGD to the Graph Isomorphism Networks (GIN) (xu2018how, ), one of the most widely applied and powerful GNNs, utilizing the MUTAG dataset (morris2020tudataset, ) as our reference graph dataset. The objective is to analyze the relationship between the input distance $GGD(G_{1},G_{2})$ and the distance between the output GIN vectors, $\|h(G_{1})-h(G_{2})\|$ for randomly selected pairs of graphs.
We employed a three-layer GIN network as described in (xu2018how, ). This network uses GIN convolutional layers to update tensors of nodes based on their neighboring nodes and then aggregates those outputs in a vector representation, followed by linear layers for classification tasks. Thus it outputs a single vector $h(G)$ for the entire graph $G$. The result is illustrated in Figure 4.
We observe a strong correlation between GGD and the output distance, as indicated by a high Pearson correlation coefficient. This finding implies the effectiveness of the proposed GGD metric for analyzing the stability of GNN models (chuang2022tree, ). To compare GGD with existing metrics, we repeat this experiment using TMD without considering node attributes (features). As shown in Figure 4, GGD demonstrates a better correlation with GIN outputs than the TMD metric across different levels. These findings indicate that when dealing with graphs without node features, GGD should be adopted for the stability analysis of graph learning models. The performance of GGD under partially missing node features is further discussed in Appendix A.10.